stuck on this one!
Solve the equation (m+8)^7/3 = -2
the choices given are: {-16} (-16,0) {-16,16} or (0)
these are my only choices

Answer by stanbon(75887)

Solve the equation (m+8)^7/3 = -2
the choices given are: {-16} (-16,0) {-16,16} or (0)
---------------
(m+8)^7/3 = -2
Raise both sides to the (3/7) power to get:
m+8= (-2)^(3/7)
m+8 = -1.345900193...
m=-9.345900193...
Cheers,
Stan H.